OpenStudy (anonymous):

Cassandra bought an antique dresser for $500. If the value of her dresser increases 6% annually, what will be the value of Cassandra's dresser at the end of 3 years to the nearest dollar?

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

A=P(1+i)^n A=500(1+0.06)^3 A=500(1.06)^3 A=500(1.19) A=595